Ammonia gas is passed into water yielding a solution of density 0.93 g/ cm^3  and containing 18.6℅ NH3 by weight . The mass of NH3 per cc of the solution is .      (1) 0.17g/cm^3  (2)  0.34g/cm^3 (3) 0.51g/cm^3 (4) 0.68 g/cm^3

Grade:12th pass

1 Answers

Arun
4 years ago

We realize that thickness = mass/volume 

Accordingly, mass = thickness * volume 

Given thickness = 0.93g/cc 

Give us a chance to ascertain mass of 1 cc of arrangement 

mass = 0.93 * 1 = 0.93g 

Presently, we realize that 1 cc of arrangement weighs 0.93g 

Given that the arrangement contains 18.6% smelling salts by weight, 1 cc of arrangement ought to have: 

18.6% of 0.93g as smelling salts 

= 0.186 * 0.93 = 0.173g 

In this manner, the mass of Ammonia per cc of arrangement = 0.173g.